Use the position equation s=-16t^2+VoT+So , where s represents the height of an object (in feet), Vo represents the inital velocity of the object(in feet per second), So represents the initial height of the object(in feet), anf T represents the time(in seconds). A projectile is fired straight upward from ground level (So=0)with a initial velocity of 224 feet per second.
a). At what instant will it be back at ground level?
b). When will the height exceed 768 feet.

Answer by Alan3354(69443)

Use the position equation s=-16t^2+VoT+So , where s represents the height of an object (in feet), Vo represents the inital velocity of the object(in feet per second), So represents the initial height of the object(in feet), anf T represents the time(in seconds). A projectile is fired straight upward from ground level (So=0)with a initial velocity of 224 feet per second.
a). At what instant will it be back at ground level?
----
-16t^2+VoT+So = 0
-16t^2 + 224t = 0
16t = 224
t = 14
=========================
b). When will the height exceed 768 feet.
-16t^2+VoT+So = 768
-16t^2 + 224t = 768
-t^2 + 14t = 48
t^2 - 14t + 48 = 0
t = 6, t = 8
=================
--> it exceeds 768 ft from 6 < t < 8